A simplified economy is specified as follows:

A. Goods market, all values C, I, G and NX values are in billions of C$:

Consumption Expenditure: C = 150 + 0.8(Y-T)Investment Expenditure: I = 1,300 - 460i Government Expenditure: G = 330Lump-sum Constant Taxes: T = 330Exports: 70 Imports: 10


B. Money market, all Md values are in billions of C$:

Interest Rate: i = 0.12 or 12%Money Demand: Md = 790 - 1,800i


Note: Please keep your answers accurate to two decimal places.

a) Given the above information, solve for the following: the equilibrium Y, the money supply M, the consumption expenditure C, and the investment expenditure I.

Find Y, M,C, and I

Please help
1
Expert's answer
2020-11-12T17:27:12-0500

Y=C+G+I+NX

Y=150 + 0.8 (Y-T)+1300 - 460i+330+(330-10)

T = 330

i=0.12

Y=150 + 0,8 (Y-T)+1300 - 460i+330+(330-10)

Y=150+0.8(Y-330)+1300-460*0.12+330+320

Y=150+0.8Y-264+1300-55.2+330+320

Y-0.8Y=1780.80

0.2Y=1780.80

Y=8904


C=150 + 0.8 (Y-T)=150+0.8(8904-330)=7009.2

I = 1,300 - 460i=1300-460*0.12=1244.8


Мd=Ms=790 - 1800i=790-1800*0.12=574